1. Below is the data provide for a particular country:
Price per computer ($) Price per cell phone ($) Quantity of computers (Unit) Quantity of cell phone (Unit)
2005 10 8 200 75
2006 14 10 250 75
2007 16 12 350 125
a) Compute nominal GDP, real GDP and GDP deflator for each year using 2005 as base year.
b) Compute the inflation rate for each year
c) Compute growth of each year.
2. Use the following data to work Problems a and b. The table lists some macroeconomic data for the
United States in 2008.
1
Expert's answer
2019-07-02T08:52:33-0400

2005 10 8 200 75

2006 14 10 250 75

2007 16 12 350 125 

a) Nominal GDP formula is: Yn = P1Q1 + P2Q2.

Yn(2005) = 10*200 + 8*75 = 2600,

Yn(2006) = 14*250 + 10*75 = 4250,

Yn(2007) = 16*350 + 12*125 = 7100.

Real GDP is calculated in base year prices, so:

Yr(2005) = 10*200 + 8*75 = 2600,

Yr(2006) = 10*250 + 8*75 = 3100,

Yr(2007) = 10*350 + 8*125 = 4500.

GDP deflator = Yn/Yr*100, so:

GDP deflator 2005 = 100,

GDP deflator 2006 = 4250/3100*100 = 137,

GDP deflator 2007 = 7100/4500*100 = 157.8.

b) Inflation rate for each year is:

In 2005 it was 0%,

in 2006 it was (137 - 100)/100*100% = 37%,

in 2007 it was (157.8 - 137)/137*100% = 15.2%.

c) Growth of each year is:

In 2006 it was (3100/2600 - 1)*100% = 19.2%,

In 2007 it was (4500/3100 - 1)*100% = 45.2%.